OpenStudy (anonymous):

i cant figure out how this it done: write ratio 2 5/8 to 8 3/4 as a fraction in simplest form.

OpenStudy (anonymous):

2 5/8 : 8 3/4 mixed number to improper fraction -- whole # times denominator plus numerator all over denominator: (2*8)+5 = 21/8 (8*4)+3 = 35/4 21/8 : 35/4 (21/8)/(35/4) (21/8)*(4/35) = 3/10

OpenStudy (anonymous):

privetek is absolutely right